OpenStudy (anonymous):

A mixture contains forty ounces of glycol and water and is ten percent glycol. If the mixture is to be strengthened to twenty-five percent, how much glycol is to be added? 4 oz 8 oz 10 oz

hero (hero):

40 + x = y 40(.10) + x = .25y

hero (hero):

Systems of equations Solve for x

hero (hero):

Put equations above in a form to be solved by elimination method 4 + x = .25y -40 - x = -y Add Fractions: -36 = -.75y Divide both sides by -1: 36 = .75y Divide both sides by .75: y = 48 x = 8 So 8 oz. But you should actually try solving these yourself rather than looking for answers.